OpenStudy (anonymous):

Is the sequence –5.4, 5.1, 15.6, 26.1, … arithmetic? If so, identify the common difference.

OpenStudy (anonymous):

Yes. To get the common difference, subtract the first term from the second, or the second term from the third, or the third term from the fourth.... and so on. Here: \[5.1-(-5.4) = 5.1+ 5.4 = 10.5 \] and you can check that if you do the same thing with the second and third term, you get the same answer, so its arithmetic.